# DACAD: Domain Adaptation Contrastive Learning for Anomaly Detection in Multivariate Time Series

### v1

@@ 3.3 Representation Layer ($\phi^{R}$)
Utilising the outputs from the representation layer $\phi^{R}$, we compute two distinct loss functions. These are the supervised mean margin contrastive loss for source domain data ($\mathcal{L}_{\text{SupCont}}$) and the self-supervised contrastive triplet loss for target domain ($\mathcal{L}_{\text{SelfCont}}$). These loss functions are critical for training our model to differentiate between normal and anomalous patterns in both source and target domains.
3.3.1 Supervised Mean Margin Contrastive Loss for Source Domain ($\mathcal{L}_{\text{SupCont}}$)
This loss aims to embed time series windows into a feature space where normal sequences are distinctively separated from anomalous ones. It utilises a triplet loss framework, comparing a base (anchor) window with both normal (positive) and anomalous (negative) windows Khosla et al. (2020). Our method diverges from traditional triplet loss by focusing on the average effect of all negatives within a batch. The formula is given by equation (1):
@@ 3.3.1 Supervised Mean Margin Contrastive Loss for Source Domain ($\mathcal{L}_{\text{SupCont}}$)
Here, $|B|$ is the batch size, and $|N|$ is the number of negative samples in the batch. The anchor $a^{s}_{i}$ is a normal time series window, and the positive pair $p^{s}_{i}$ is another randomly selected normal window. Negative pairs $n^{s}_{j}$ are either true anomalous windows or synthetically created anomalies through the anomaly injector module applied on the anchor $a^{s}_{i}$. This loss function uses true anomaly labels to enhance the separation between normal and anomalous behaviours by at least the margin $m$. It includes both genuine and injected anomalies as negatives, balancing ground truth and potential anomaly variations. The supervised mean margin contrastive loss offers several advantages for TAD: supervised learning (uses labels for better anomaly distinction), comprehensive margin (applies a margin over average distance to negatives), and flexible negative sampling (incorporates a mix of real and injected anomalies, enhancing robustness against diverse anomalous patterns).
3.3.2 Self-supervised Contrastive Triplet Loss for Target Domain ($\mathcal{L}_{\text{SelfCont}}$)
For the target domain, we employ a self-supervised contrastive approach using triplet loss Schroff et al. (2015), designed to ensure that the anchor window is closer to a positive window than to a negative window by a specified margin. The self-supervised triplet loss formula is shown in equation (2):
@@ 3.4 Discriminator Component ($\phi^{D}$)
Designed to differentiate between features extracted from the source and target domains, the discriminator employs adversarial training techniques similar to those found in Generative Adversarial Networks (GANs) Creswell et al. (2018). Discriminator $\phi^{D}$ is trained to differentiate between source and target domain features, while $\phi^{R}$ is conditioned to produce domain-invariant features.
A crucial element is the Gradient Reversal Layer Ganin and Lempitsky (2015), which functions normally during forward passes but reverses gradient signs during backpropagation. This setup enhances the training of $\phi^{D}$ and simultaneously adjusts $\phi^{R}$ to produce features that challenge $\phi^{D}$.
@@
3.5 Centre-based Entropy Classifier ($\phi^{CL}$)
Extending the DeepSVDD Ruff et al. (2018) designed for anomaly detection, the CEC ($\phi^{CL}$) in DACAD is proposed as an effective anomaly detection classifier in the source domain. It assigns anomaly scores to time series windows, using labelled data from the source domain $S$ for training and applying the classifier to target domain data $T$ during inference. It is centred around a Multi-Layer Perceptron (MLP) with a unique “centre“ parameter crucial for classification.
The classifier operates by spatially separating transformed time series window representations ($\phi^{R}(w_{i}^{s})$) in the feature space relative to a predefined “centre“ $c$. The MLP’s aim is to draw normal sample representations closer to $c$ and push anomalous ones further away. This spatial reconfiguration is quantified using a distance metric, forming the basis for anomaly scoring. A sample closer to $c$ is considered more normal, and vice versa. The effectiveness of this spatial adjustment is measured using a loss function based on BCE, expressed in equation (4):
$\mathcal{L}_{\text{Cls}}=-\frac{1}{|S|}\sum_{i=1}^{|S|}\left[y_{i}\cdot\log(\|g(w_{i}^{s})-c\|_{2}^{2})+(1-y_{i})\cdot\log(1-\|g(w_{i}^{s})-c\|_{2}^{2})\right]$ (4)
$where~{}g(w)=\phi^{CL}(\phi^{R}(w))$
@@ 3.5 Centre-based Entropy Classifier ($\phi^{CL}$)
The loss function is designed to minimise the distance between normal samples and $c$ while maximising it for anomalous samples. These distances are directly used as anomaly scores, offering a clear method for anomaly detection.
3.6 Overall Loss in DACAD
@@
The overall loss function $\mathcal{L}_{\text{DACAD}}$ is what the model seeks to optimise during the training process. By fine-tuning these weights ($\alpha$, $\beta$, $\gamma$, and $\lambda$), the model can effectively balance the significance of each loss component. This balance is crucial as it allows the model to cater to specific task requirements.
3.7 DACAD’s Inference
In the inference phase of the DACAD model, the primary objective is to identify anomalies in the target domain $T$. This is accomplished for each window $w^{t}$ in $T$. The anomaly detection is based on the concept of spatial separation in the feature space, as established during the training phase. The anomaly score for each window is derived from the spatial distance between the classifier’s output and a predefined centre $c$ in the feature space. This score is calculated as the squared Euclidean distance, as shown in Equation 6:
@@ 3.7 DACAD’s Inference
Where $\phi^{CL}(\phi^{R}(w^{t}))$ denotes the feature representation of the window $w^{t}$ after processing by the representation layer $\phi^{R}$ and the classifier layer $\phi^{CL}$. The distance measured here quantifies how much each window in the target domain deviates from the “normal“ pattern, as defined by the centre $c$.
@@
In practical terms, the anomaly score can be thresholded to classify windows as either normal or anomalous. By effectively utilising these anomaly scores, the DACAD model provides a robust mechanism for identifying anomalous patterns in unlabelled target domain data.
4 Experiments
The aim of this section is to provide a comprehensive evaluation of DACAD, covering the experimental setups (Section 4.2) and results (Section 4.3 and 4.4), to clearly understand its performance and capabilities in different contexts.
4.1 Datasets
We evaluate the performance of the proposed model and make comparisons of the results across the four datasets, including the three most commonly used real benchmark datasets for TAD and the dataset used for time series domain adaptation. The datasets are summarised in Table 1.
Table 1: Statistics of the benchmark datasets used.
@@ 4.1 Datasets
MSL Hundman et al. (2018) 27 55 58,317 73,729 10.72%
SMAP Hundman et al. (2018) 55 25 140,825 444,035 13.13%
SMD Su et al. (2019) 28 38 708,405 708,420 4.16%
Boiler Cai et al. (2021) 3 274 277,152 277,152 15.00%
4.2 Experimental Setup
@@
In our study, we evaluate SOTA TAD models including OmniAnomaly Su et al. (2019), TS2Vec Yue et al. (2022), THOC Shen et al. (2020), and DCdetector Yang et al. (2023) and SOTA UDA models that support multivariate time series classification including VRADA Purushotham et al. (2016) and CLUDA Ozyurt et al. (2023) on benchmark datasets previously mentioned in section 1 using their source code and best hyper-parameters as they stated to ensure a fair evaluation. The hyper-parameters used in our implementation are as follows: DACAD consists of a 3-layer TCN architecture with three different channel sizes [128, 256, 512] to capture temporal dependencies. The dimension of the representation is 1024. We use the same hyper-parameters across all datasets to evaluate DACAD: window size ($WS$) = 100, margin $m$ = 1, and we run our model for 20 epochs.
@@
Table 2: F1, AUPR, and AUROC results for various models on multivariate time series benchmark datasets (SMD, MSL, SMAP). The most optimal evaluation results are displayed in bold, while the second best ones are indicated by underline.
@@
Anomaly Detection OmniAnom F1 0.397 0.413 0.419 0.421 0.254 0.266 0.246 0.242 0.320 0.326 0.331 0.318
@@ 4.3 Baselines Comparison
Table 2 provides a comprehensive comparison of various models’ performance on multivariate time series benchmark datasets, using F1 score, AUPR (Area Under the Precision-Recall Curve), and AUROC (Area Under the Receiver Operating Characteristic Curve). Despite Point Adjustment (PA) popularity in recent years, we do not use PA when calculating these metrics due to Kim et al.’s findings that its application leads to an overestimation of a TAD model’s capability and can bias results in favour of methods that produce extreme anomaly scores. Instead, we use conventional performance metrics for anomaly detection.
@@
Benchmark datasets like SMD contain multiple time series that cannot be merged due to missing timestamps, making them unsuitable for averaging their F1 scores. The F1 score is a non-additive metric combining precision and recall. To address this, we compute individual confusion matrices for each time series. These matrices are then aggregated into a collective confusion matrix for the entire dataset. From this aggregated matrix, we calculate the overall precision, recall, and F1 score, ensuring an accurate and undistorted representation of the dataset’s F1 score. For datasets with multiple time series, we present the mean and standard deviation of AUPR and AUROC for each series.
@@
DACAD emerges as the top performer in Table 2, consistently achieving the best results across all scenarios and metrics, as highlighted in bold. This suggests its robustness and adaptability in handling normal and anomalous representations of time series for anomaly detection. VRADA and CLUDA models often rank as the second best, with their results underlined in several instances. Other models like OmniAnomaly, THOC, TS2Vec, and DCdetector demonstrate more uniform performance across various metrics but generally do not reach the top performance levels seen in DACAD, VRADA, or CLUDA. Their consistent but lower performance could make them suitable for applications where top-tier accuracy is less critical.

### v2

@@ III-C Representation Layer ($\phi^{R}$)
Utilizing the outputs from the representation layer $\phi^{R}$, we compute two distinct loss functions. These are the supervised mean margin contrastive loss for source domain data ($\mathcal{L}_{\text{SupCont}}$) and the self-supervised contrastive triplet loss for target domain ($\mathcal{L}_{\text{SelfCont}}$). These loss functions are critical for training our model to differentiate between normal and anomalous patterns in both source and target domains.
III-C1 Supervised Mean Margin Contrastive Loss for Source Domain ($\mathcal{L}_{\text{SupCont}}$)
This loss aims to embed time series windows into a feature space where normal sequences are distinctively separated from anomalous ones. It utilizes a triplet loss framework, comparing a base (anchor) window with both normal (positive) and anomalous (negative) windows [59]. Our method diverges from traditional triplet loss by focusing on the average effect of all negatives within a batch. The formula is given by equation (1):
@@ III-C1 Supervised Mean Margin Contrastive Loss for Source Domain ($\mathcal{L}_{\text{SupCont}}$)
Here, $|B|$ is the batch size, and $|N|$ is the number of negative samples in the batch. The anchor $a^{s}_{i}$ is a normal time series window, and the positive pair $p^{s}_{i}$ is another randomly selected normal window. Negative pairs $n^{s}_{j}$ are either true anomalous windows or synthetically created anomalies through the anomaly injector module applied on the anchor $a^{s}_{i}$. This loss function uses true anomaly labels to enhance the separation between normal and anomalous behaviors by at least the margin $m$. It includes both genuine and injected anomalies as negatives, balancing ground truth and potential anomaly variations. The supervised mean margin contrastive loss offers several advantages for TSAD: supervised learning (uses labels for better anomaly distinction), comprehensive margin (applies a margin over average distance to negatives), and flexible negative sampling (incorporates a mix of real and injected anomalies, enhancing robustness against diverse anomalous patterns).
III-C2 Self-supervised Contrastive Triplet Loss for Target Domain ($\mathcal{L}_{\text{SelfCont}}$)
For the target domain, we employ a self-supervised contrastive approach using triplet loss [58], designed to ensure that the anchor window is closer to a positive window than to a negative window by a specified margin. The self-supervised triplet loss formula is shown in equation (2):
@@ III-D Discriminator Component ($\phi^{D}$)
Designed to differentiate between features extracted from the source and target domains, the discriminator employs adversarial training techniques similar to those found in Generative Adversarial Networks (GANs) [62]. Discriminator $\phi^{D}$ is trained to differentiate between source and target domain features, while $\phi^{R}$ is conditioned to produce domain-invariant features.
A crucial element is the Gradient Reversal Layer [63], which functions normally during forward passes but reverses gradient signs during backpropagation. This setup enhances the training of $\phi^{D}$ and simultaneously adjusts $\phi^{R}$ to produce features that challenge $\phi^{D}$.
@@
III-E Centre-based Entropy Classifier ($\phi^{CL}$)
Extending the DeepSVDD [14] designed for anomaly detection, the CEC ($\phi^{CL}$) in DACAD is proposed as an effective anomaly detection classifier in the source domain. It assigns anomaly scores to time series windows, using labeled data from the source domain $S$ for training and applying the classifier to target domain data $T$ during inference. It is centered around a Multi-Layer Perceptron (MLP) with a unique “centre“ parameter crucial for classification.
The classifier operates by spatially separating transformed time series window representations ($\phi^{R}(w_{i}^{s})$) in the feature space relative to a predefined “centre“ $c$. The MLP aims to draw normal sample representations closer to $c$ and push anomalous ones further away. This spatial reconfiguration is quantified using a distance metric, forming the basis for anomaly scoring. A sample closer to $c$ is considered more normal, and vice versa. The effectiveness of this spatial adjustment is measured using a loss function based on BCE, expressed in equation (III-E):
$\mathcal{L}_{\text{Cls}}=$ $-\frac{1}{|S|}\sum_{i=1}^{|S|}\left[y_{i}\cdot\log(\|g(w_{i}^{s})-c\|_{2}^{2})\right.$
$\left.+(1-y_{i})\cdot\log(1-\|g(w_{i}^{s})-c\|_{2}^{2})\right]$
where $\text{where }g(w)=\phi^{CL}(\phi^{R}(w))$ (4)
@@ III-E Centre-based Entropy Classifier ($\phi^{CL}$)
The loss function is designed to minimize the distance between normal samples and $c$ while maximizing it for anomalous samples. These distances are directly used as anomaly scores, offering a clear method for anomaly detection.
III-F Overall Loss in DACAD
@@
The overall loss function $\mathcal{L}_{\text{DACAD}}$ is what the model seeks to optimize during the training process. By fine-tuning these weights ($\alpha$, $\beta$, $\gamma$, and $\lambda$), the model can effectively balance the significance of each loss component. This balance is crucial as it allows the model to cater to specific task requirements.
III-G DACAD’s Inference
In the inference phase of the DACAD model, the primary objective is to identify anomalies in the target domain $T$. This is accomplished for each window $w^{t}$ in $T$. The anomaly detection is based on the concept of spatial separation in the feature space, as established during the training phase. The anomaly score for each window is derived from the spatial distance between the classifier’s output and a predefined centre $c$ in the feature space. This score is calculated as the squared Euclidean distance, as shown in Equation (6):
@@ III-G DACAD’s Inference
Where $\phi^{CL}(\phi^{R}(w^{t}))$ denotes the feature representation of the window $w^{t}$ after processing by the representation layer $\phi^{R}$ and the classifier layer $\phi^{CL}$. The distance measured here quantifies how much each window in the target domain deviates from the “normal” pattern, as defined by the centre $c$.
@@
In practical terms, the anomaly score can be thresholded to classify windows as either normal or anomalous. By effectively utilizing these anomaly scores, the DACAD model provides a robust mechanism for identifying anomalous patterns in unlabelled target domain data.
IV Experiments
This section provides a comprehensive evaluation of DACAD, covering the experimental setups (Section V-A) and results (Section V-B and V-E), to clearly understand its performance and capabilities in different contexts.
TABLE I: Statistics of the benchmark datasets used.
@@ IV Experiments
MSL [37] 27 55 58,317 73,729 10.72%
SMAP [37] 55 25 140,825 444,035 13.13%
SMD [44] 28 38 708,405 708,420 4.16%
Boiler [8] 3 274 277,152 277,152 15.00%
IV-A Datasets
We evaluate the proposed model and make comparisons across the four datasets, including the three most commonly used real benchmark datasets for TSAD and the dataset used for time series domain adaptation. The datasets are summarised in Table I.
Mars Science Laboratory (MSL) and Soil Moisture Active Passive (SMAP) [^1] https://www.kaggle.com/datasets/patrickfleith/nasa-anomaly-detection-dataset-smap-msl datasets [37] are real-world datasets collected from NASA spacecraft. These datasets contain anomaly information derived from reports of incident anomalies for a spacecraft monitoring system. MSL and SMAP comprise 27 and 55 datasets, respectively, and each is equipped with a predefined train/test split, where, unlike other datasets, their training set is unlabeled.
Server Machine Dataset (SMD) [^2] https://github.com/NetManAIOps/OmniAnomaly/tree/master/ServerMachineDataset [44] is gathered from 28 servers, incorporating 38 sensors, over a span of 10 days. During this period, normal data was observed within the initial 5 days, while anomalies were sporadically injected during the subsequent 5 days. The dataset is also equipped with a predefined train/test split, where the training data is unlabeled.
Boiler Fault Detection Dataset. [^3] https://github.com/DMIRLAB-Group/SASA-pytorch/tree/main/datasets/Boiler [8] The Boiler dataset includes sensor information from three separate boilers, with each boiler representing an individual domain. The objective of the learning process is to identify the malfunctioning blowdown valve in each boiler. Obtaining samples of faults is challenging due to their scarcity in the mechanical system. Therefore, it’s crucial to make use of both labeled source data and unlabeled target data to enhance the model’s ability to generalize.
V Baselines
Below, we will provide an enhanced description of the UDA models for time series classification and anomaly detection. Additionally, we provide a description of the five prominent and state-of-the-art TSAD models that were used for comparison with DACAD. We have selected the models from different categories of TSAD, namely, unsupervised reconstruction-based (OmniAnomaly [44] and AnomalyTransformer [39]) model, unsupervised forecasting-based (THOC [47]), and self-supervised contrastive learning (TS2Vec [36] and DCdetector [53]) TSAD models.
AE-MLP [64] is an anomaly detection model that uses autoencoders for nonlinear dimensionality reduction. It encodes data points into a lower-dimensional space and reconstructs them, highlighting deviations to detect anomalies.
AE-LSTM [65] is an anomaly detection model using an LSTM-based encoder-decoder for multi-sensor data. It detects anomalies by reconstruction error of time series sequences.
RDC [^4] https://github.com/syorami/DDC-transfer-learning [66] is built upon deep domain confusion to enhance domain adaptation by maximizing domain invariance. It aligns feature distributions through a domain confusion objective, making learned representations indistinguishable across domains and improving predictive model performance on unseen domains.
VRADA [^5] https://github.com/floft/vrada [7] combines deep domain confusion [66] with variational recurrent adversarial deep domain adaptation [7] optimizing source domain label prediction, MMD, and domain discrimination using latent representations from the LSTM encoder. Concurrently, AE-LSTM’s reconstruction objective is utilized for anomaly detection.
SASA [^6] https://github.com/DMIRLAB-Group/SASA [8] improves time series classification using adaptation across domains by identifying and aligning key sparse patterns. It uses a self-attention layer with LSTM units to find optimal global context windows for source and target domains and aligns them using Maximum Mean Discrepancy (MMD).
CLUDA [^7] https://github.com/oezyurty/CLUDA [9] is a novel framework developed for UDA in time series data and evaluated on time series classification. This framework utilizes a contrastive learning approach to learn domain-invariant semantics within multivariate time series data, aiming to preserve label information relevant to prediction tasks. CLUDA is distinctive as it is the first UDA framework designed to learn contextual representations of time series data, maintaining the integrity of label information. Its effectiveness is demonstrated through evaluations of various time series classification datasets.
ContextDA [^8] There is no implementation available for this model, so we rely on the results claimed in the paper. [13] is a sophisticated approach for detecting anomalies in time series data. It formulates context sampling as a Markov decision process and employs deep reinforcement learning to optimize the domain adaptation process. This model is designed to generate domain-invariant features for better anomaly detection across various domains. It has shown promise in transferring knowledge between similar or entirely different domains.
OmniAnomaly [^9] https://github.com/smallcowbaby/OmniAnomaly [44] is a model operating on an unsupervised basis, employing a Variational Autoencoder (VAE) to handle multivariate time series data. It identifies anomalies by evaluating the reconstruction likelihood of specific data windows.
THOC [^10] We utilized the authors’ shared implementation, as it is not publicly available. [47] utilizes a multi-layer dilated recurrent neural network (RNN) alongside skip connections in order to handle contextual information effectively. It adopts a temporal hierarchical one-class network approach for detecting anomalies.
TS2Vec [^11] https://github.com/yuezhihan/ts2vec [36] is an unsupervised model that is capable of learning multiple contextual representations of MTS and UTS semantically at various levels. This model employs contrastive learning in a hierarchical way, which provides a contextual representation. A method within TS2Vec has been proposed for application in TSAD.
AnomalyTransformer [^12] https://github.com/thuml/Anomaly-Transformer [39] detects anomalies in time series data by using a novel ”Anomaly-Attention” mechanism to compute association discrepancies, enhancing the distinguishability between adjacent normal and abnormal points through a minimax strategy.
DCdetector [^13] https://github.com/DAMO-DI-ML/KDD2023-DCdetector [53] is distinctive for its use of a dual attention asymmetric design combined with contrastive learning. Unlike traditional models, DCdetector does not rely on reconstruction loss for training. Instead, it utilizes pure contrastive loss to guide the learning process. This approach enables the model to learn a permutation invariant representation of time series anomalies, offering superior discrimination abilities compared to other methods.
V-A Experimental Setup
@@
In our study, we evaluate several SOTA TSAD models including OmniAnomaly [44], TS2Vec [36], THOC [47], and DCdetector [53] and SOTA UDA models that support multivariate time series classification including VRADA [7] and CLUDA [9] on benchmark datasets previously mentioned in section I using their source code and best hyper-parameters as they stated to ensure a fair evaluation. The hyper-parameters used in our implementation are as follows: DACAD consists of a 3-layer TCN architecture with three different channel sizes [128, 256, 512] to capture temporal dependencies. The dimension of the representation is 1024. We use the same hyper-parameters across all datasets to evaluate DACAD: window size ($WS$) = 100, margin $m$ = 1, and we run our model for 20 epochs.
@@
TABLE II: F1, AUPR, and AUROC results for various models on multivariate time series benchmark datasets (SMD, MSL, SMAP). The most optimal evaluation results are displayed in bold, while the second best ones are indicated by underline.
@@
Aff-F1 0.816$\pm$0.103 0.823$\pm$0.105 0.813$\pm$0.101 0.791$\pm$0.124 0.760$\pm$0.090 0.743$\pm$0.082 0.748$\pm$0.093 0.735$\pm$0.088 0.738$\pm$0.099 0.753$\pm$0.114 0.757$\pm$0.123 0.757$\pm$0.128
Anomaly Detection Omni [44] F1 0.397 0.413 0.419 0.421 0.254 0.266 0.246 0.242 0.320 0.326 0.331 0.318
@@ V-B Baselines Comparison
Table II provides a comprehensive comparison of the performance of the different models on multivariate time series benchmark datasets. The performance measures include the F1 score, AUPR (Area Under the Precision-Recall Curve), and AUROC (Area Under the Receiver Operating Characteristic Curve). Despite Point Adjustment (PA) popularity in recent years, we do not use PA when calculating these metrics due to Siwon Kim’s findings  [67] that its application leads to an overestimation of a TSAD model’s capability and can bias results in favor of methods that produce extreme anomaly scores. Instead, we use conventional performance metrics for anomaly detection.
@@
Benchmark datasets like SMD contain multiple time series that cannot be merged due to missing timestamps, making them unsuitable for averaging their F1 scores. The F1 score is a non-additive metric combining precision and recall. To address this, we compute individual confusion matrices for each time series. These matrices are then aggregated into a collective confusion matrix for the entire dataset. From this aggregated matrix, we calculate the overall precision, recall, and F1 score, ensuring an accurate and undistorted representation of the dataset’s F1 score. Additionally, we report the Affiliation F1 (Aff-F1) [68], calculated from Aff-Precision (Aff-Pre) and Aff-Recall (Aff-Rec), to provide a comprehensive evaluation focused on early detection. For datasets with multiple time series, we present the mean and standard deviation of Aff-F1, AUPR, and AUROC for each series.
@@
It is evident from Table II that DACAD is the best performer, consistently achieving the best results across all scenarios and metrics, as highlighted in bold. This suggests its robustness and adaptability to handling normal and anomalous representations of time series for anomaly detection. VRADA and CLUDA models often rank as the second best, with their results underlined in several instances. Other models like OmniAnomaly, THOC, TS2Vec, AnomalyTransformer, and DCdetector demonstrate more uniform performance across various metrics but generally do not reach the top performance levels seen in DACAD, VRADA, or CLUDA. Their consistent but lower performance could make them suitable for applications where top-tier accuracy is less critical.
The addition of the Affiliation F1 (Aff-F1) metric further reinforces the strength of DACAD. A high value of the affiliation metric indicates a strong overlap or alignment between the detected anomalies and the true anomalies in a time series. DACAD’s superior Aff-F1 scores across all datasets underline its efficacy in accurately identifying anomalies, highlighting its potential as a solution in time series anomaly detection.
TABLE III: Macro F1/AUROC results on SMD and Boiler dataset. The most optimal evaluation results are displayed in bold, while the second best ones are indicated by underline.
Data Models (Macro F1/AUROC)
src $\mapsto$ trg AE-MLP AE-LSTM RDC VRADA SASA CLUDA ContexTDA DACAD
SMD 1-1 $\mapsto$ 1-2 0.72 / 0.83 0.74 / 0.90 0.74 / 0.89 0.74 / 0.91 0.59 / 0.63 0.75 / 0.84 0.75 / 0.91 0.80 / 0.84
1-1 $\mapsto$ 1-3 0.57 / 0.70 0.49 / 0.41 0.57 / 0.72 0.49 / 0.41 0.61 / 0.90 0.67 / 0.77 0.57 / 0.75 0.71 / 0.84
1-1 $\mapsto$ 1-4 0.55 / 0.74 0.54 / 0.41 0.54 / 0.75 0.54 / 0.41 0.55 / 0.75 0.56 / 0.74 0.59 / 0.76 0.75 / 0.83
1-1 $\mapsto$ 1-5 0.54 / 0.79 0.55 / 0.84 0.56 / 0.85 0.55 / 0.79 0.65 / 0.87 0.68 / 0.87 0.66 / 0.87 0.82 / 0.96
1-1 $\mapsto$ 1-6 0.71 / 0.88 0.71 / 0.91 0.71 / 0.87 0.71 / 0.91 0.44 / 0.84 0.71 / 0.77 0.73 / 0.84 0.89 / 0.92
1-1 $\mapsto$ 1-7 0.48 / 0.55 0.48 / 0.50 0.49 / 0.54 0.48 / 0.50 0.31 / 0.57 0.50 / 0.62 0.51 / 0.53 0.88 / 0.91
1-1 $\mapsto$ 1-8 0.55 / 0.57 0.53 / 0.70 0.55 / 0.58 0.54 / 0.56 0.52 / 0.56 0.56 / 0.61 0.58 / 0.58 0.79 / 0.72
Boiler 1 $\mapsto$ 2 0.44 / 0.64 0.43 / 0.48 0.43 / 0.54 0.43 / 0.48 0.53 / 0.88 0.49 / 0.52 0.50 / 0.59 0.61 / 0.66
1 $\mapsto$ 3 0.40 / 0.28 0.40 / 0.11 0.43 / 0.49 0.42 / 0.18 0.41 / 0.48 0.47 / 0.79 0.50 / 0.67 0.74 / 0.81
2 $\mapsto$ 1 0.39 / 0.18 0.40 / 0.21 0.40 / 0.36 0.40 / 0.15 0.53 / 0.90 0.47 / 0.49 0.51 / 0.66 0.58 / 0.65
2 $\mapsto$ 3 0.40 / 0.38 0.40 / 0.20 0.45 / 0.39 0.42 / 0.21 0.41 / 0.49 0.48 / 0.50 0.50 / 0.69 0.59 / 0.63
3 $\mapsto$ 1 0.39 / 0.20 0.40 / 0.16 0.39 / 0.31 0.40 / 0.15 0.48 / 0.67 0.47 / 0.77 0.51 / 0.67 0.72 / 0.84
3 $\mapsto$ 2 0.48 / 0.54 0.49 / 0.48 0.49 / 0.55 0.49 / 0.48 0.46 / 0.31 0.49 / 0.57 0.50 / 0.57 0.58 / 0.68
V-C UDA Comparison
Table III, adapted from [13], now includes results of CLUDA and our model in the two last columns. This table provides a comprehensive comparison of the performances of various models on the SMD and Boiler datasets, as measured by Macro F1 and AUROC scores.